# Bribe & Fork: Cheap PCN Bribing Attacks via Forking Threat

### v1

@@ Appendix 0.B Notation Summary
$tx_{u},tx_{1},tx_{2},tx_{b},tx_{p_{1}},tx_{p_{2}}$ Transactions used to construct Bribe & Fork
(see the paragraph ”Types of transaction sets.”).
$txs_{u},txs_{1},txs_{2},txs_{p_{1}},txs_{p_{2}}$ Transaction sets that include transactions
used to construct Bribe & Fork (see the paragraph
”Types of transaction sets.”).
@@
Table 1: Notation summary
Appendix 0.C Analysis of Bribe & Fork
In this section we formally analyze Bribe & Fork where a bribe transaction $tx_{2}$ is published, large enough to bribe a chosen miner with the highest mining power, yet significantly smaller than the value required to directly persuade all miners to skip mining the transaction $txs_{1}$. The selected miner is then asked to threaten others with a fork if they add the unwanted transaction $txs_{1}$ to the blockchain. To make these threats credible, we implemented the self-penalty mechanism (see Section 2.2)
About the proofs. In the proofs presented in this section, we aim to find a dominating strategy for every player $N_{i}$ in a given state $\mathcal{S}$ and a round $r$. Whenever needed, we use the mathematical induction technique to show that some choice of strategy is optimal for a sequence of rounds. Usually, the base case is the last round of the game and the induction step proves that if a given strategy is dominating in a round $k+1$, then it is also a dominating strategy in a round $k$.
@@ Appendix 0.C Analysis of Bribe & Fork
0.C.1 Transaction Order in a Single Chain
Although, due to the definition of the player’s utility function, the bribing transaction ($txs_{2}$) may encourage the players not to skip mining transactions with high rewards during the dispute period (e.g. $txs_{1}$), we first show that once timelock is over and $txs_{1}$ was not mined, the players will follow the default strategy to mine transactions with highest rewards first.
@@ 0.C.1 Transaction Order in a Single Chain
0.C.2 Decisions of an Individual Miner are Consistent
[…]
@@ 0.C.2 Decisions of an Individual Miner are Consistent
0.C.3 Only a High-Cost Reward May Encourage Miners to Fork
@@
[…]
@@
By Theorem 4, the player $N_{j}$ that mined $txs_{1}$, will continue to mine the chain with $txs_{1}$ in the next two rounds. Furthermore, as already mentioned, any miner $N_{i}$ that decides to continue the chain with $txs_{1}$, can profit at least $\lambda_{i}(f+B)$ in each round till the end of the game, by simply mining $txs_{u}$. From the above it follows that.
@@
$\boldsymbol{\mathbb{E}}_{i}(\sigma_{2},\mathcal{S}_{0,0}^{i,0,1},r+1)\leq\mathsf{reward}_{i}(\mathcal{S})+\max\{(f_{2}+B)(1-\lambda_{j})+(M-1)\lambda_{i}(f+B)$
@@
0.C.4 Without a High-Cost Reward, All Players Mine $txs_{1}$
As we already observed, once $txs_{1}$ is mined, it will not be forked unless the bribing fee is sufficient. We will show that for a sufficiently large number of rounds $T$, all of the players will mine transaction set $txs_{1}$ in the first round. A similar result was introduced in [26], but we prove that this result still holds in the game with forks.
[…]
@@ 0.C.4 Without a High-Cost Reward, All Players Mine $txs_{1}$
In the game where none of the miners decides to create the transaction set $txs_{p_{1}}$, miners may choose to mine $txs_{u}$ and $txs_{1}$ in all rounds, or $txs_{2}$ only after round $T$. Now, since the game contains only one pair of mutually exclusive transactions $txs_{1},txs_{2}$ with $\mathsf{reward}(txs_{2})>\mathsf{reward}(txs_{1})$, then by Conjecture 1 players can start to fork only when $txs_{1}$ appears on the blockchain. What is more, since $f_{2}-f<10^{-2}(f+B)$, by Theorem 4, whenever some player with $\lambda_{j}>10^{-2}$ successfully mines $txs_{1}$ in a chain of length $\leq T$, none of the players will decide to fork his block.
@@
Now, $\boldsymbol{\mathbb{E}}_{i}(\sigma_{2},\mathcal{S}_{T},T)<\boldsymbol{\mathbb{E}}_{i}(\sigma_{1},\mathcal{S}_{T},T)$ only if (*$(*)f_{2}-f\geq\frac{f_{1}-f}{\lambda_{i}}$. This implies that miners with collective mining power at least $Y$ will prefer to mine $txs_{1}$ in this round.
@@
Given a game with forks $\Gamma_{F}(\mathcal{S},r)$ with $r<T$, where none of the miners decides to create $txs_{p_{1}}$ before the first round (or $txs_{u}$ is mined in the first round), and the bribing fee is not too high, i.e. $f_{2}-f<10^{-2}(f+B)$ and given $Y=\sum_{j=i:\lambda_{j}>0.01,f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}}^{|N|}\lambda_{j}$; $T-r+1,Y$ are big enough, such that $(1-1.01(1-Y)^{T-r+1})>0$, every miner with $\lambda_{j}>0.01$ will decide to mine $f_{1}$ in this round.
0.C.5 Discouraging Miners to Mine $txs_{1}$

### v2

@@ 3.4 Notation Summary
4 The Attack
In Section 2.4, we gave a comment on the design choices made for our model, and in Section 3, we gave a detailed description of our model describing the mining game with forks. Next, we analyze our attack where a bribing transaction $tx_{2}$ is created that is high enough to bribe a selected miner with the highest mining power (but the value is significantly smaller than the value needed to directly convince all miners in the game to skip mining the transaction $txs_{1}$) and he is asked to threaten other players that he will fork them, once they put the unwanted transaction $txs_{1}$ on the blockchain. For this, we introduced the self-penalty mechanism (Section 2.3) that makes the threats of the selected miner credible.
We prove that the above strategy is effective in a sequence of steps. In Section 4.1 we show that once the dispute period is over (miners mine on a chain of length $\geq T$), then all the players will mine the transactions with higher fees first. Next, in the Section 4.2 we show that without the self-penalty mechanism, each miner that already mined some block, will continue to mine this block even when some other player starts a fork of this block, and in the Section 4.3 we show that without the self-penalty mechanism, majority of the miners will not decide to fork any transaction in the game. These two results let us prove in Section 4.4 that in the game with forks where the self-penalty mechanism is not used, the miners will choose to mine $txs_{1}$ in the first round (similarly as in the game without forks defined in [zeta]). On the other hand in Section 4.5, we define a penalty that forces the selected miner to fork $txs_{1}$, once it appears on blockchain, and poses a threat strong enough, that all the miners avoid mining $txs_{1}$ before the chain length $T$ is reached. In the Section 4.6 we show that, in summary, it is more profitable for all players to use the self-penalty mechanism and mine $txs_{2}$, than to mine the $txs_{1}$ directly.
About the proofs. In the proofs presented in this section, we often find in a given state $\mathcal{S}$ and a round $r$ a dominating strategy for every player $N_{i}$. Whenever needed, we use the mathematical induction technique to show that some choice of strategy is optimal for a sequence of rounds. Usually, the base case is the last round of the game and the induction step proves that if a given strategy is dominating in a round $k+1$, then it is also a dominating strategy in a round $k$.
@@ 4 The Attack
4.1 Transaction Order in a Single Chain
Although, due to the definition of the player’s utility function, the bribing transaction ($txs_{2}$) may encourage the players not to skip mining transactions with high rewards during the dispute period (e.g. $txs_{1}$), we show that once timelock is over and $txs_{1}$ was not mined, the players will follow the default strategy to mine transactions with highest rewards first.
@@ 4.1 Transaction Order in a Single Chain
4.2 Decisions of an Individual Miner are Consistent
In this section we show that without a high-cost reward $f_{2}$, once someone is successful with mining $txs_{1}$, the miner will continue mining this chain, as it might be too costly for the miner to lose the block reward that he already mined. As $txs_{1}$, $txs_{2}$ is the only pair of conflicting transactions in the game whenever $txs_{p_{1}}$ was not created, it follows from Conjecture 1 that the forks may occur only when one miner successfully mines $txs_{1}$, and the other player wants to profit from mining $txs_{2}$. Thus, in the following, we study the behaviour of the players whenever one of the players decides to mine $txs_{1}$.
Theorem 4.1
Assuming subgame $\Gamma_{F}(\mathcal{S},r)$ in a state $\mathcal{S}$ with a single chain of length $r\leq T$, formed until round $r$ where player $N_{j}$ mined $txs_{1}$ in the last round, and $txs_{p_{1}}$ is not on the chain, then the player $N_{j}$ will $\mathsf{continue}$ to mine this chain unless $f_{2}-f\geq f_{1}+B$, even when other miners decide to fork the chain with $txs_{1}$ and $\mathsf{continue}$ mining the new subchain created during the fork.
@@ 4.2 Decisions of an Individual Miner are Consistent
4.3 Only a High-Cost Reward May Encourage Miners to Fork
@@
Theorem 4.2
Let $\Gamma_{F}(\mathcal{S},r)$ be a subgame, where $txs_{p_{1}}$ was not created, $r\leq T$, and the state $\mathcal{S}$ contains only a single chain of length $r$ consisting of $r-1$ unrelated transaction sets $txs_{u}$ and one (just mined) $txs_{1}$ (mutually exclusive with $txs_{2}$ with $\mathsf{reward}(txs_{2})>\mathsf{reward}(txs_{1})>\mathsf{reward}(txs_{u})$) mined by some miner $N_{j}$. In the game $txs_{p_{1}}$ was not created. Other miners will not fork $txs_{1}$, unless $f_{2}-f\geq\lambda_{j}(f+B)$, where $\lambda_{j}$ is the mining power of the miner $N_{j}$.
@@
By Theorem 4.1, the player $N_{j}$ that mined $txs_{1}$, will continue to mine the chain with $txs_{1}$ in the next two rounds. Furthermore, as already mentioned, any miner $N_{i}$ that decides to continue the chain with $txs_{1}$, can profit at least $\lambda_{i}(f+B)$ in each round till the end of the game, by simply mining $txs_{u}$. From the above it follows that.
@@
$\boldsymbol{\mathbb{E}}_{i}(\sigma_{2},\mathcal{S}_{0,0}^{i,0,1},r+1)\leq\mathsf{reward}_{i}{}_{i}(\mathcal{S})+\max\{(f_{2}+B)(1-\lambda_{j})+(M-1)\lambda_{i}(f+B)$
@@
4.4 Without a High-Cost Reward, All Players Mine $txs_{1}$
As we already observed, once $txs_{1}$ is mined, it will not be forked unless the bribing fee is sufficient. We will show that for a sufficiently large number of rounds $T$, all of the players will mine transaction set $txs_{1}$ in the first round. A similar result was introduced in [timelocked], but we prove that this result still holds in the game with forks.
Theorem 4.3
Given a game with forks $\Gamma_{F}(\mathcal{S},1)$ where none of the miners decides to create $txs_{p_{1}}$ before the first round, and the bribing fee is not too high, i.e. $f_{2}-f<10^{-2}(f+B)$, $f_{1}>f$, and given $Y=\sum_{j=i:\lambda_{j}>0.01,f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}}^{+\infty}\lambda_{j}$; $T,Y$ are big enough, such that $(1-1.01(1-Y)^{T})>0$, every miner with $\lambda_{j}>0.01$ will decide to mine $f_{1}$ in the first round.
@@ 4.4 Without a High-Cost Reward, All Players Mine $txs_{1}$
In the game where none of the miners decides to create the transaction set $txs_{p_{1}}$, miners may choose to mine $txs_{u}$ and $txs_{1}$ in all rounds, or $txs_{2}$ only after round $T$. Now, since the game contains only one pair of mutually exclusive transactions $txs_{1},txs_{2}$ with $\mathsf{reward}(txs_{2})>\mathsf{reward}(txs_{1})$, then by Conjecture 1 players can start to fork only when $txs_{1}$ appears on the blockchain. What is more, since $f_{2}-f<10^{-2}(f+B)$, by Theorem 4.2, whenever some player with $\lambda_{j}>10^{-2}$ successfully mines $txs_{1}$ in a chain of length $\leq T$, none of the players will decide to fork his block.
@@
Now, $\boldsymbol{\mathbb{E}}_{i}(\sigma_{2},\mathcal{S}_{T},T)<\boldsymbol{\mathbb{E}}_{i}(\sigma_{1},\mathcal{S}_{T},T)$ only if $(*)f_{2}-f\geq\frac{f_{1}-f}{\lambda_{i}}$. This implies that miners with collective mining power at least $Y$ will prefer to mine $txs_{1}$ in this round.
@@
Given a game with forks $\Gamma_{F}(\mathcal{S},r)$ with $r<T$, where none of the miners decides to create $txs_{p_{1}}$ before the first round (or $txs_{u}$ is mined in the first round), and the bribing fee is not too high, i.e. $f_{2}-f<10^{-2}(f+B)$ and given $Y=\sum_{j=i:\lambda_{j}>0.01,f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}}^{+\infty}\lambda_{j}$; $T-r+1,Y$ are big enough, such that $(1-1.01(1-Y)^{T-r+1})>0$, every miner with $\lambda_{j}>0.01$ will decide to mine $f_{1}$ in this round.
4.5 Discouraging Miners to Mine $txs_{1}$
In the previous sections, we have shown that it is rather expensive to force the players not to mine $txs_{1}$ in the first round, even when the players can fork this transaction. In this section, we leverage the self-penalty mechanism introduced in Section 2.3. The proof is inductive, and its base case starts in round $T$. For each round, we first show that the miner $N_{s}$ with the highest mining power $\lambda_{s}$ will not mine $txs_{1}$, as we assume that $f_{2}-f>\frac{f_{1}-f}{\lambda_{s}}$. Next, given a sufficiently large penalty $P>\lambda_{s}(f+B)$, we show that the selected player $N_{s}$ will fork the transaction $txs_{1}$, once it appears on the blockchain, even though it poses a risk of loosing the block reward. Finally, we show that in this round all players other than the player $N_{s}$ are afraid to mine $txs_{1}$, when the self-penalty transaction is on the chain.
Theorem 4.4
Assuming that $f_{2}-f>\frac{f_{1}-f}{\lambda_{s}}$, $\frac{f+B}{f_{1}+B}>1-\lambda_{s}^{2}$ and $txs_{p_{1}}$ was mined in the first round with $P>\lambda_{s}(f+B)$ by the player $N_{s}$ with mining power $\lambda_{s}$, then none of the miners will decide to mine $txs_{1}$ in rounds $2,\ldots,T$.
Because of the space constraints, proof of the Theorem 4.4 was moved to the Appendix 0.A.
4.6 Encouraging the Strongest Miner to Use the Penalty Mechanism
Finally, we observe the benefit that comes from using the penalty mechanism. First for the miner with the strongest mining power $\lambda_{s}$, we observe that using the self-penalty mechanism and threatening others to mine the transaction $txs_{1}$, once it appears on the blockchain is beneficial for him whenever $f_{2}-f>\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}+\overline{f}$. Next, for any miner with a smaller mining power, we show that merely the fact that he is threatened to mine $txs_{1}$ can force them to skip mining this transaction.
Theorem 4.5
In the game with forks $\Gamma_{F}$ that starts with an empty state $\mathcal{S}$, whenever $f_{2}-f>\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}+\overline{f}$, $\frac{f+B}{f_{1}+B}>1-\lambda_{s}^{2}$, $f_{1}>f$, $f_{p_{2}}>f$, $\lambda_{min}>0.05^{T/2}$, $f_{2}-f<10^{-2}(f+B)$ and given $Y=\sum_{j=i:\lambda_{j}>0.01,f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}}^{+\infty}\lambda_{j}$, it holds that $(1-1.01(1-(1-Y)^{T/2})>0$, the dominating strategy for all players in the game $\Gamma_{F}$ is is to mine $txs_{p_{1}}$ with $P>\lambda_{s}(f+B)$ created in the first round by the strongest player $N_{s}$ with the mining power $\lambda_{s}$, then mine $txs_{u}$ until round $T$, then $txs_{2}$, $txs_{p_{2}}$, and $txs_{u}$ until the end.
Proof
By Theorems 4.3 and  4.4, utility of the player $N_{s}$ that chooses to create $txs_{p_{1}},txs_{p_{2}}$ and mine $txs_{p_{1}}$ in the first round [^4] In the analysis we omit the strategy where the player $N_{s}$ creates $txs_{p_{1}},txs_{p_{2}}$ and does not decide to mine $txs_{p_{1}}$. (strategy $\sigma_{p}$) is at least:
where $F^{\prime}_{1}=(\lambda_{s}(T-1)(f+B)+\lambda_{s}(R-T)(f+B)),$
Recall that we assume that all players with mining power less than $1\%$ have collective power at most $5\%$. As the players with mining power more than $0.01$ will prefer to mine $txs_{1}$ in the first place when $txs_{p_{1}}$ is not created, the utility of the player $N_{s}$ that does not decide to create $txs_{p_{1}}$ (strategy $\sigma_{1}$) is at most (by Lemma 3 and the Theorem 4.4):
On other hand the players may first decide to mine either $txs_{u}$ or $txs_{1}$ in the first round (strategy $\sigma_{1*}$). In the worst case scenario the block with $txs_{1}$ is not forked. What is more, whenever $1-\lambda_{p_{1}}$ miners decide to mine $txs_{u}$ in the first round, then all miners with mining power more than $0.01$ will make an attempt to mine $txs_{1}$. In conclusion, by Lemma 3 and the Theorem 4.4:
Which holds for any $f_{p_{1}}\geq f_{1}$ and $\lambda_{i}>0.05^{T/2}$.
5 Experimental Analysis
We begin with the analysis of the fee structures that can be found in the real-world data. Next, we study the distribution of hash power. Finally, we present a sample evaluation of the cost of the bribing attack with forks.
5.1 Fee Structure
We analyze the real-world data on fees earned by miners on Bitcoin. In particular, we explore the Bitcoin historical data from 2022 ([btcexplorer]). In this period, the block reward was the sum of the block reward base, 6.25 BTC, and the fees for including transactions. The weekly average sum of the block fees ranged between 0.05 and 0.225 BTC as shown in Figure 3. The average fee of a single transaction fluctuated proportionally between $0.00004$ and $0.00016$ BTC as shown in Figure 4. We can thus fairly assume that any block typically contains on average $1500$ transactions, the sum of fees of an average block of transactions is around $0.15$ BTC and the average fee of a single transaction is approximately $10^{-4}$ BTC.
Figure 3: Average amount of fees aggregated in a block for each week in 2022 ([btcexplorer]).
Figure 4: Average transaction fee of a single transaction for each week in 2022 ([btcexplorer]).
5.2 Hash Power Analysis
Device Hashrate (Hashes/second)
Nvidia GeForce RTX 4090 $2.75\cdot 10^{8}$
Nvidia GeForce RTX 3090 $9.4\cdot 10^{7}$
Bitmain Antminer S19 XP Hyd $2.55\cdot 10^{14}$
Baikal BK-G28 $2.8\cdot 10^{10}$
Table 1: Hash rates of commonly used Bitcoin mining devices [asicminer].
The weakest miner. In the previous work [zeta], the authors assumed that the fraction of the mining power held by the weakest miners is approximately $10^{-3}$. The data depicted in the Table 1 and in the Figure $2(b)$ shows that it is a highly unrealistic assumption for Bitcoin. A very strong GPU which is often used by individual miners for mining has hashrate of around $3\cdot 10^{8}$ (Table 1), whereas the collective mining power in Bitcoin can be estimated to $2\cdot 10^{20}$ (Figure $2(b)$). In conclusion, we cannot expect that every individual miner has more than $10^{-12}$ probability of mining a block.
Strong miners. On the other side of the spectrum, top mining pools continuously maintain over $20\%$ of total hashpower. To calculate the share of the hashpower of a big mining pool in a given period, we count the number of the blocks mined by it. Then we divide this number by the total number of mined blocks in the period. Figure 5(a) and Figure 6 in the Appendix show that, in the most popular blockchains, most of the time, one of the miners (mining pools) has the collective power higher than $20\%$.
Pool Number of mined blocks Percentage of total mined blocks
Foundry USA $11851$ $22.281\%$
AntPool $8670$ $16.301\%$
F2Pool $7848$ $14.755\%$
Binance Pool $6164$ $11.589\%$
ViaBTC $5162$ $9.705\%$
Poolin $4438$ $8.344\%$
BTC.com $2252$ $4.234\%$
SlushPool $2142$ $4.027\%$
Luxor $1362$ $2.561\%$
SBI Crypto $963$ $1.811\%$
Braiins Pool $715$ $1.344\%$
unknown $698$ $1.312\%$
MARA Pool $307$ $0.577\%$
Terra Pool $149$ $0.280\%$
ULTIMUS POOL $147$ $0.276\%$
KuCoinPool $101$ $0.190\%$
EMCDPool $80$ $0.150\%$
PEGA Pool $76$ $0.143\%$
Table 2: The scale of mining hashrates of the most significant mining pools and independent miners of Bitcoin in 2022. Data sources: [btcexplorer].
Figure 5: (a) Collective power of top mining pools in 2022 for selected Blockchains, monthly data [btcexplorer][ltc][etc]. (b) Collective number of hashes per second calculated by all miners in 2022 [$10^{20}$ hashes/second], weekly data [hashpower].
5.3 Example Evaluation of The Bribing Attack With Forks
In this section, we present a sample evaluation that visualizes the improvement that our bound $f_{2}-f>\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}+\overline{f}$ makes in comparison to the older result from [zeta], i.e., $f_{2}-f\geq\frac{f_{1}-f}{\lambda_{min}}$. We also show that there exists a value of $f_{2}$ which does not violate the other condition of the theorem, i.e., $\sum_{j=i:\lambda_{j}>0.01,f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}}^{+\infty}\lambda_{j}\geq Y$.
For example, let us assume that $f_{1}-f\approx\overline{f}$, and set $T>110$. Now, since $\lambda_{min}$ can be fairly estimated to be $\lambda_{min}<10^{-12}$, we can see that the attack without forking threats could cost in practice around $10^{12}\cdot\overline{f}$. On the other hand, the new bound requires only $f_{2}-f>\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}+\overline{f}$, for $\lambda_{s}\approx 0.2$, this costs around $f_{2}-f\geq 21\overline{f}$. The only condition left is that for some miner with $\lambda_{j}>0.01$, the following condition must hold $f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}$, but the data shows that miners that control approximately $1.5\%-2\%$ of the total mining power usually exist, thus for a miner with mining power $1\%<\lambda_{j}<2\%$ it holds that $f_{2}-f<50\cdot\overline{f}$. In summary, if we take any $f_{2}$ that is larger than $f$ by $21$ up to $50$ times, then the default strategy for all miners is to wait for the bribing transaction.
6 Related Work and Countermeasures
In the landscape of constructing financially stable systems on blockchain [miller2017zerocollateral, financiallystable2], our work falls into the class of incentive manipulation attacks. Specifically, to the best of our knowledge, we are the first to combine feather forking attacks [perspectives] with timelock bribing attacks on the payment channel networks and to achieve a bribing cost that is approximately only constant times larger than the cost of an average transaction fee.
Timelock bribing was originally introduced by Khabbazian et al. [timelocked]. Later, Avarikioti et al. [zeta] applied timelock bribing attacks in payment channel networks, such as the Lightning Network and Duplex Micropayment Channels, and proposed countermeasures. Our work extends [zeta], modifying the timelock bribing attack for payment channels to facilitate a miner bribing strategy that incorporates feather forking.
The bribing strategies for the payment channels are similar in their nature to the bribing strategies for the HTLC mechanism. In [temporary], the authors introduced a way to bribe the HTLC mechanism, leveraging the power of smart contracts and feather forking. The cost of the attack in this work is however proportional to the sum of the fees ($\gtrsim\sum_{i=1}^{T}f\cdot\lambda_{max}$) of all transactions before the deadline $T$. Perhaps the most notable approach to mitigating timelocked bribing was presented in MAD-HTLC [madhtlc]. They show that the HTLC mechanism is susceptible to the bribing attacks similar to the attack presented in [zeta] (achieving exactly the same attack cost $\approx\frac{f_{1}-f}{\lambda_{min}}$, as the one presented by Avarikioti et al. [zeta], independent of the number of rounds before the deadline $T$). The authors of [madhtlc] assume however that the minimal fraction of mining power controlled by a single user is at least $0.01$, which is a big constraint compared to our work. Additionally in this work, a modification of the Hashed Time-Locked Contracts (HTLC) was proposed to allow miners to collect the coins locked in the HTLC in case a party misbehaves (again similar to the mechanism proposed in [zeta]). He-HTLC [helium] pointed out the counterbribing vulnerabilities in MAD-HTLC. Their work proposes a bribe that is bigger than the amount of a collateral used to secure assets in the construction proposed in MAD-HTLC, but requires the proactivity of the miners. The value of collateral may be however much bigger than the average fees, as it is supposed to be collected back whenever the players play honestly and it it hard to compare the size of the bribing fee to other results. Additionally, [helium] and  [rapidash] suggested another secure modifications against bribes that leverage the power of the proactivity of the miners. In contrast to these works, our focus is not on designing countermeasures against timelocked bribing attacks. Instead, we demostrate how employing feather forking can make timelocked bribing attacks very cheap for the attacker, therefore highlighting the need for robust mitigating strategies.
To defend against our proposed attack, one may leverage the mitigation technique for timelocked bribing on the Lighting Network presented by Avarikioti et al. [zeta]. In the context of our model, the countermeasure assures that announcing $txs_{2}$ comes with announcing some secret that can be used by anyone to claim the money before time $T$. This means that if the $txs_{2}$ is announced on the mining pool before time $T$, all of the money to be collected only after time $T$ can be immediately claimed by some other party. We leave it as a fact without proof that the same countermeasure mechanism still works even in the model that assumes forks. Intuitively, the “strong” miner in our analysis does not profit from waiting for the bribing transaction if it is not announced, preventing the creation of the self-penalty transaction. Contrarily, if the transaction is announced and the secret is revealed, any (winning) miner could claim the reward money.
7 Conclusions and Future Work
In conclusion, our work reexamines the vulnerability of PCNs to bribing attacks and introduces a modified attack leveraging the threat of forking. We introduce a formal model of a mining game with forking extending the conditionally timelocked games introduced by Avarikioti et al. [zeta]. In particular, in our extended model, miners not only choose which transactions to mine in each round but also decide whether to continue existing chains or initiate forks. In this model, we demonstrate that the cost of the bribing attack can be significantly reduced compared to the previous analysis. In more detail, it can be reduced from $\frac{f_{1}-f}{\lambda_{min}}$ to approximately $\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}$, where $\overline{f}$ represents the cost of an average fee for a single transaction and $\lambda_{s}$ denotes the reduction factor compared to significantly smaller $\lambda_{min}$ calculated in previous work [zeta]. To validate our findings, we empirically analyze the historical data of real-world blockchain implementations. This analysis confirms that staging a bribing attack on a PCN is significantly less costly than considered previously.
The results of our study have implications for the design and implementation of PCNs, as well as for the broader applications of timelocked contracts, e.g., atomic swaps. Our findings underscore the need for proactive measures to mitigate the risk of bribing attacks. Possible avenues for future research include exploring whether the penalty mechanism implementation described in Section 2.4 can be implemented without the honest majority assumption or whether our results still hold in the presence of more general abandon rules.